The first overtone of an open pipe has frequency n. The first overtone of a closed pipe of the same length will have frequency

Grade:12

1 Answers

Arun
5 years ago
Dear studentfor open organ pipe fundamental frequency is f1 = v/2l.for closed organ pipe fundamental frequency is f2 = v/4l.f1/f2 = 2Hence Frequency of closed organ pipes = n/2